40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 28 20:42:53 2020, Koji, Update, General, Vent Prep 10x
    Reply  Thu Oct 29 12:43:35 2020, gautam, Update, General, Vent Prep preVentAlignment.pngPRMIcarr_PRMalignment.pngALS_ool.pdflowPowerMC.png
       Reply  Thu Oct 29 19:30:56 2020, gautam, Update, General, Vent complete ventValveStatus.pngvent_Oct2020.jpgventValveStatus_overnight.pnglowPowerIMC_afterVent.png
          Reply  Mon Nov 2 12:44:25 2020, gautam, Update, General, ETMY and OMC chamber heavy doors off CB84FD62-C48B-414F-8825-50C9DE0B47CF.jpeg
             Reply  Mon Nov 2 16:46:06 2020, gautam, Update, General, Plan for OMC chamber IMG_2318.JPGIMG_2332.JPG
                Reply  Mon Nov 2 17:13:19 2020, Koji, Update, General, Plan for OMC chamber 
                   Reply  Mon Nov 2 17:32:05 2020, gautam, Update, General, Plan for OMC chamber IMG_2317.JPG
Message ID: 15651     Entry time: Thu Oct 29 12:43:35 2020     In reply to: 15649     Reply to this: 15652
Author: gautam 
Type: Update 
Category: General 
Subject: Vent Prep 
  1. Oplev HeNe at ETMY was replaced, see here for my earlier discussion on this.
    • I thought this is a good idea since we want the Oplev as a coarse reference and it'd not be ideal if this HeNe dies during the time the optic is out of the chamber.
    • New HeNe had 2.8mW of power output as measured with Ophir power meter. This is in line with what is expected from these Lumentum heads.
    • I labelled the head with the power output and today's date, re-aligned the Oplev reflection onto its QPD. 
    • After this work, the Y-arm could be locked without the huge angular fluctuations that was visible earlier, 👍 .
  2. GTRY anomaly
    • I actually judged that there is no anomaly.
    • The GTRY CDS indicator is actually quite useless - the ADC saturates at ~3500 cts (and not 32768 as you would expect from a 16 bit ADC but that's the well-known whitening filter saturation problem). This should be fixed, but this is a task for later.
    • I measured with a DMM the voltage when the TEM00 is locked to the cavity and GTRY is 0.3 (the nominal value these days), the DC voltage was ~5.6 V. The prompt reflection from the ETM registers ~6.5 V DC. So the mode-matching isn't stellar, but this is again a known issue, and can be fixed later.
  3. Other pre-vent checks
    • The Oplevs had drifted significantly, I re-centered ITMs, ETMs and BS after aligning the arm cavities and green beams in the POX/POY lock state. See Attachment #1.
    • I locked the PRMI on carrier and used this config to re-center the PRM Oplev, see Attachment #2.
    • No further action was taken regarding SRM oplev.
    • I checked the ALS noise, see Attachment #3. The X arm ALS has excess noise at ~100 Hz that certainly wasn't there previosuly - sigh. There is nothing I can find about any changes made at EX in the elog.
    • Updated the "DriftMon" values, though I guess we don't even really use this anymore these days?
    • Re-relieved the IMC WFS offsets.
    • Cut the input power to the IMC from 1.007W to 100.1 mW (both numbers measured with Ophir power meter).
    • Replaced the 10% beamsplitter in the MCREFL path on the AS table with a Y1 HR mirror. Note that there is no beam on the IMC WFS in this configuration.
    • Was able to lock the IMC on low power to a TEM00 mode - need to set up the low power autolocker. The IMC autolocker is now set to the low power settings, and I've tested it locks a couple of times. Attachment #5 shows the low power lock in StripTool.
    • Walked around and looked at all the bellows - the jam nuts are up against their stops, and I can't move them with my hands, so I think that's okay.

If everything else looks good, I'll start letting the dry N2 into the main volume after lunch.

Quote:

Now the green transmissions are visible by the green PDs. Attachment 9 shows the trans and ref of each green beams with and without locking to TEM00. The questionable green TRY was ~0.3. If we compare this with the histrical data (Attachment 10), it is about 1/4 of the value in the past. It's not too crazy but still quite low.

BTW, nice video! @ Koji, How difficult was it to edit it into this form? 

Attachment 1: preVentAlignment.png  283 kB  Uploaded Thu Oct 29 14:33:13 2020  | Hide | Hide all
preVentAlignment.png
Attachment 2: PRMIcarr_PRMalignment.png  235 kB  Uploaded Thu Oct 29 14:33:18 2020  | Hide | Hide all
PRMIcarr_PRMalignment.png
Attachment 3: ALS_ool.pdf  518 kB  Uploaded Thu Oct 29 14:35:31 2020  | Hide | Hide all
ALS_ool.pdf
Attachment 4: lowPowerMC.png  55 kB  Uploaded Thu Oct 29 15:59:57 2020  | Hide | Hide all
lowPowerMC.png
ELOG V3.1.3-